# Building symmetries into data-driven manifold dynamics models for complex flows

### v1

@@ IV.2.3 Long-time predictions
Another important quantity to consider is the ability of the models to capture the energy balance of the system. In Fig. 18 we show the joint PDF of $I$ and $D$ for the DNS and the same models. Again, the model trained in the fundamental space closely matches the joint PDF of the true data. The phase-aligned model both underestimates the energy associated with the high probability RPOs, and overestimates the energy associated with the low probability high power input and dissipation events.
@@ V Conclusion
Symmetries appear naturally in dynamical systems and we show here that correctly accounting for them dramatically improves the performance of data-driven models for time evolution on an invariant manifold. In this work, we introduce a method that we call symmetry charting and apply it to Kolmogorov flow in a chaotic regime of $\text{Re}=14.4$. This symmetry charting method factors out symmetries so we can train ROMs in a fundamental chart and ensure equivariant trajectories – we are essentially learning coordinates and dynamics on one region (chart) of the invariant manifold for the long-time dynamics rather than having to learn these for the entire manifold. To do so, we first factor out the symmetries by identifying a set of indicators that differentiate the set of discrete symmetries for the system. Here, Fourier coefficients serve this purpose; we found that the signs of specific Fourier modes could uniquely identify all discrete symmetry operations. We then use IRMAE-WD, an autoencoder architecture that tends to drive the rank of the latent space covariance of the data to a minimum, to find a low-dimensional representation of the data. This method overcomes the need to sweep over latent space dimensions. We observe that factoring out symmetries improves the MSE of the reconstruction as well as the dimension estimates and robustness to IRMAE-WD hyperparameters. When considering the original data (i.e. no symmetries factored out) and the phase-aligned data (only continuous symmetry factored out) the MSE is higher and the range of dimension estimates is wider. We also note that the dimension estimate range in the fundamental chart agrees well with the dimension found in our previous work for the same system [10]. We then train a NODE with latent space of dimension $d_{h}=10$ for both phase-aligned and fundamental space data. This dimension is the upper bound of the dimension estimate obtained for the fundamental chart data. The resulting models using symmetry charting to map to the fundamental space accurately reconstructed the DNS at both short and long times. In contrast, the phase-aligned model quickly landed on an unphysical stable RPO leading to poor reconstruction of the dynamics.
@@
The approach described here is essentially a version of the “CANDyMan” (Charts and Atlases for Nonlinear Data-driven Dynamics on Manifolds) approach described in [15, 16]. As shown on those studies, learning charts and atlases to develop local manifold representations and dynamical models can improve performance for systems with complex dynamics. There, however, the charts were found by clustering; here that step is bypassed by using a priori knowledge, about the system, namely its symmetries. The methodology presented in this work can be applied to other dynamical systems with rich symmetries by identifying the corresponding indicators. Future directions for symmetry charting and CANDyMan include applications such as control in systems with symmetry (cf. [20]) as well as development of hierarchical methods where the fundamental chart identified by symmetry can be further subdivided using a cluster algorithm or a chart autoencoder [37].

### v2

@@ IV.2.3 Long-time predictions
Another important quantity to consider is the ability of the models to capture the energy balance of the system. In Fig. 18 we show the joint PDF of $I$ and $D$ for the DNS and the same models. Again, the model trained in the fundamental domain closely matches the joint PDF of the true data. The phase-aligned model both underestimates the energy associated with the high probability RPOs, and overestimates the energy associated with the low probability high power input and dissipation events.
@@ V Conclusion
Symmetries appear naturally in dynamical systems and we show here that correctly accounting for them dramatically improves the performance of data-driven models for time evolution on an invariant manifold. In this work, we introduce a method that we call symmetry charting and apply it to Kolmogorov flow in a chaotic regime of $\text{Re}=14.4$. This symmetry charting method factors out symmetries so we can train ROMs in a fundamental chart and ensure equivariant trajectories – we are essentially learning coordinates and dynamics on one region (chart) of the invariant manifold for the long-time dynamics rather than having to learn these for the entire manifold. To do so, we first factor out the symmetries by identifying a set of indicators that differentiate the set of discrete symmetries for the system. Here, Fourier coefficients serve this purpose; we found that the signs of specific Fourier modes could uniquely identify all discrete symmetry operations. We then use IRMAE-WD, an autoencoder architecture that tends to drive the rank of the latent space covariance of the data to a minimum, to find a low-dimensional representation of the data. This method overcomes the need to sweep over latent space dimensions. We observe that factoring out symmetries improves the MSE of the reconstruction as well as the dimension estimates and robustness to IRMAE-WD hyperparameters. When considering the original data (i.e. no symmetries factored out) and the phase-aligned data (only continuous symmetry factored out) the MSE is higher and the range of dimension estimates is wider. We also note that the dimension estimate range in the fundamental chart agrees well with the dimension found in our previous work for the same system [10]. We then train a NODE with latent space of dimension $d_{h}=10$ for both phase-aligned and fundamental domain data. This dimension is the upper bound of the dimension estimate obtained for the fundamental chart data. The resulting models using symmetry charting to map to the fundamental domain accurately reconstructed the DNS at both short and long times. In contrast, the phase-aligned model quickly landed on an unphysical stable RPO leading to poor reconstruction of the dynamics.
@@
The approach described here is essentially a version of the “CANDyMan” (Charts and Atlases for Nonlinear Data-driven Dynamics on Manifolds) approach described in [15, 16]. As shown in those studies, learning charts and atlases to develop local manifold representations and dynamical models can improve performance for systems with complex dynamics. There, however, the charts were found by clustering; here that step is bypassed by using a priori knowledge, about the system, namely its symmetries. The methodology presented in this work can be applied to other dynamical systems with rich symmetries by identifying the corresponding indicators. Future directions for symmetry charting and CANDyMan include applications such as control in systems with symmetry (cf. [20]) as well as development of hierarchical methods where the fundamental chart identified by symmetry can be further subdivided using a cluster algorithm or a chart autoencoder [47].
@@ Acknowledgements.
Appendix A Symmetry operations in Fourier space
In this appendix we show the derivations to obtain the actions of the symmetry operations on the Fourier coefficients as shown in Equations 7 through 9. We start with the shift (in $y$)-reflect (in $x$) symmetry
$\mathscr{S}:[u,v,\omega](x,y)\rightarrow[-u,v,-\omega]\left(-x,y+\frac{\pi}{n}\right).$ (24)
Throughout this appendix we will focus on the vorticity representation $\omega(x,y)$, which we can write in terms of its Fourier coefficients $\hat{\omega}(k_{x},k_{y})$ as
$\omega(x,y)=\sum_{k_{x}}\sum_{k_{y}}\hat{\omega}(k_{x},k_{y})e^{i(k_{x}x+k_{y}y)}.$ (25)
The symmetry operations act term-by-term on the Fourier series. Applying the shift (in $y$)-reflect (in $x$) symmetry to each term yields
$\mathscr{S}:\hat{\omega}(k_{x},k_{y})e^{i(k_{x}x+k_{y}y)}\rightarrow-\hat{\omega}(k_{x},k_{y})e^{-ik_{x}x+ik_{y}y+ik_{y}\frac{\pi}{n}}.$ (26)
We can simplify the equation to get
$\mathscr{S}:\hat{\omega}(k_{x},k_{y})\rightarrow-\hat{\omega}(-k_{x},k_{y})e^{ik_{y}\frac{\pi}{n}}.$ (27)
For the rotation through $\pi$,
$\mathscr{R}:\omega(x,y)\rightarrow\omega(-x,-y),$ (28)
we can also write in Fourier representation
$\mathscr{R}:\hat{\omega}(k_{x},k_{y})e^{i(k_{x}x+k_{y}y)}\rightarrow\hat{\omega}(k_{x},k_{y})e^{-i(k_{x}x+ik_{y}y)}.$ (29)
Simplifying this equation we get
$\mathscr{R}:\hat{\omega}(k_{x},k_{y})\rightarrow\hat{\omega}(-k_{x},-k_{y}).$ (30)
Since $\hat{\omega}(k_{x},k_{y})$ and $\hat{\omega}(-k_{x},-k_{y})$ are complex conjugates, we can write this as
$\mathscr{R}:\hat{\omega}(k_{x},k_{y})\rightarrow\bar{\hat{\omega}}(k_{x},k_{y}).$ (31)
For the continuous translation in $x$,
$\mathscr{T}_{l}:\omega(x,y)\rightarrow\omega(x+l,y),$ (32)
we can write in Fourier representation
$\mathscr{T}_{l}:\hat{\omega}(k_{x},k_{y})e^{i(k_{x}x+k_{y}y)}\rightarrow\hat{\omega}(k_{x},k_{y})e^{ik_{x}x+ik_{y}y+ik_{x}l)}.$ (33)
Simplifying this equation we get
$\mathscr{T}_{l}:\hat{\omega}(k_{x},k_{y})\rightarrow\hat{\omega}(k_{x},k_{y})e^{ik_{x}l}.$ (34)
Appendix B Symmetry operations and indicators for plane Couette and Poiseuille flows
Here we show how to select a set of indicators for plane Couette flow (PCF) and plane Poiseuille flow (PPF) with periodic boundary conditions in $x$ and $z$ and no-slip boundary conditions at $y=-1$ and $y=1$. For PCF, the Navier-Stokes equations and boundary conditions are invariant to two discrete symmetry operations ($\sigma_{C1}$, $\sigma_{C2}$) and a continuous translation $\mathscr{T}_{C}(l_{x},l_{z})$ [48]:
$\sigma_{C1}:[u,v,w](x,y,z)\rightarrow[u,v,-w](x,y,-z),$ (35)
$\sigma_{C2}:[u,v,w](x,y,z)\rightarrow[-u,-v,w](-x,-y,z),$ (36)
$\mathscr{T}_{C}(l_{x},l_{z}):[u,v,w](x,y,z)\rightarrow[u,v,w](x+l_{x},y,z+l_{z}).$ (37)
The operation $\sigma_{C1}$ corresponds to reflecting the velocity field across the $x-z$ plane, and $\sigma_{C2}$ to rotating it around the $x$ axis. Similarly, for PPF, the Navier-Stokes equations and boundary conditions are invariant to two discrete symmetries ($\sigma_{P1}$, $\sigma_{P2}$) and a continuous translation $\mathscr{T}_{P}(l_{x},l_{z})$ [49]:
$\sigma_{P1}:[u,v,w](x,y,z)\rightarrow[u,v,-w](x,y,-z),$ (38)
$\sigma_{P2}:[u,v,w](x,y,z)\rightarrow[u,-v,w](x,-y,z),$ (39)
$\mathscr{T}_{P}(l_{x},l_{z}):[u,v,w](x,y,z)\rightarrow[u,v,w](x+l_{x},y,z+l_{z}).$ (40)
Note that $\sigma_{C1}=\sigma_{P1}$ and $\mathscr{T}_{C}(l_{x},l_{z})=\mathscr{T}_{P}(l_{x},l_{z})$, while $\sigma_{C2}$ and $\sigma_{P2}$ differ in whether $u$ changes sign (Poiseuille) or not (Couette) under rotation.
The translation symmetries in $x$ and $z$ can be treated just as we did for Kolmogorov flow, using the first Fourier-mode method of slices in $x$ and $z$. Since there are two discrete symmetry operations, each data point lies in one of four discrete subspaces, and our goal is to find criteria that uniquely identify which one.
For PCF and PPF, it is useful to consider the action of these symmetry operations on the Fourier-Chebyshev representation of the flow, given by
$\mathbf{u}(x,y,z)=\sum_{k_{x}}\sum_{k_{y}}\sum_{k_{z}}\hat{\mathbf{u}}(k_{x},k_{y},k_{z})T_{k_{y}}(y)e^{i(k_{x}x+k_{z}z)},$ (41)
where $\mathbf{u}=[u,v,w]$, $\hat{\mathbf{u}}=[\hat{u},\hat{v},\hat{w}]$, and $T_{k_{y}}(y)$ are Chebyshev polynomials. Under this transformation, the symmetry operations are
$\sigma_{C1}:[\hat{u},\hat{v},\hat{w}](k_{x},k_{y},k_{z})\rightarrow[\hat{u},\hat{v},-\hat{w}](k_{x},k_{y},-k_{z}),$ (42)
$\sigma_{C2}:[\hat{u},\hat{v},\hat{w}](k_{x},k_{y},k_{z})\rightarrow(-1)^{k_{y}}[-\hat{u},-\hat{v},\hat{w}](-k_{x},k_{y},k_{z}),$ (43)
$\sigma_{P2}:[\hat{u},\hat{v},\hat{w}](k_{x},k_{y},k_{z})\rightarrow(-1)^{k_{y}}[\hat{u},-\hat{v},\hat{w}](k_{x},k_{y},k_{z}),$ (44)
$\mathscr{T}_{C}(l_{x},l_{z}):[\hat{u},\hat{v},\hat{w}](k_{x},k_{y},k_{z})\rightarrow[\hat{u},\hat{v},\hat{w}](k_{x},k_{y},k_{z})e^{i(k_{x}l_{x}+k_{z}l_{z})}.$ (45)
Now, we can consider specific values of $(k_{x},k_{y},k_{z})$ to select our criteria. Basing the criteria on $k_{x}=0$ and $k_{z}=0$ coefficients decouples them from phase-alignment with the method-of-slices. Additionally, we can use the fact that $\sigma_{C2}$ and $\sigma_{P2}$ modify $\hat{v}(0,k_{y},0)$ and $\hat{w}(0,k_{y},0)$ in the same way to identify the same criteria for both PCF and PPF. We do this by considering the effect of the discrete symmetry operations on $\hat{v}(0,2,0)$:
$\sigma_{C1},\sigma_{P1}:\hat{v}(0,2,0)\rightarrow\hat{v}(0,2,0),$ (46)
$\sigma_{C2}:\hat{v}(0,2,0)\rightarrow-\hat{v}(0,2,0),$ (47)
$\sigma_{P2}:\hat{v}(0,2,0)\rightarrow-\hat{v}(0,2,0),$ (48)
and on $\hat{w}(0,2,0)$:
$\sigma_{C1},\sigma_{P1}:\hat{w}(0,2,0)\rightarrow-\hat{w}(0,2,0),$ (49)
$\sigma_{C2}:\hat{w}(0,2,0)\rightarrow\hat{w}(0,2,0),$ (50)
$\sigma_{P2}:\hat{w}(0,2,0)\rightarrow\hat{w}(0,2,0).$ (51)
Observe that $\sigma_{C1}(=\sigma_{P1})$ changes the sign of $\hat{w}(0,2,0)$ but not $\hat{v}(0,2,0)$, so we use the former as a criterion for this operation. For $\sigma_{C2}$ and $\sigma_{P2}$, the opposite occurs, so we use the sign of $\hat{v}(0,2,0)$ as the criterion for those operators. In summary, we can define criteria, and corresponding integer indicators $\mathcal{I}$, based on the quadrant in ($\hat{v}(0,2,0)$,$\hat{w}(0,2,0)$) in which the data lies, as shown in Table 3.
Table 3: Indicators $\mathcal{I}$, corresponding operations required to map to $\mathcal{I}=0$, the fundamental domain, and signs of the relevant Fourier-Chebyshev coefficients, for PCF and PPF.
 $\begin{array}[]{cccc}\hline\cr\hline\cr\mathcal{I}&\text{\; \; \; \; Discrete operation }&\textrm{sgn}(\hat{v}(0,2,0))&\textrm{sgn}(\hat{w}(0,2,0))\\ \hline\cr 0&\text{Identity}&+&+\\ 1&\sigma_{C1},\sigma_{P1}&+&-\\ 2&\sigma_{C2},\sigma_{P2}&-&+\\ 3&\sigma_{C1}\sigma_{C2},\sigma_{P1}\sigma_{P2}&-&-\\ \hline\cr\hline\cr\end{array}$